# Alleviating Hallucinations in Large Vision-Language Models through Hallucination-Induced Optimization

### v1

@@ 6.1 Experimental Settings
Benchmarks. We evaluate HIO on three benchmarks including (1) quantitative metrics POPE Li et al. (2023b) on MSCOCO Lin et al. (2014) dataset. The Polling-based Object Probing Evaluation Li et al. (2023b) offers a streamlined approach to assessing object hallucination. In this benchmark, LVLMs are queried about the existence of specific objects in a given image. (2)CHAIR Rohrbach et al. (2018), Caption Hallucination Assessment with Image Relevance, is a specialized tool designed to evaluate the occurrence of object hallucination in image captioning tasks. (3) general-purposed Multimodal Large Language Model Evaluation (MME) Fu et al. (2023) benchmark, which provides an extensive benchmark designed to evaluate LVLMs across multiple dimensions, including ten perception-related subtasks and four cognition-focused ones.
LVLM Baselines We evaluate the effectiveness of our HIO on three state-of-the-art LVLMs. Concretely, we apply our HIO to LLaVA-1.5, which employ Vicuna 7B as language decoder Liu et al. (2023c). For a more convincing comparison, we report the averaged results as well as the standard deviation over $5$ runs on POPE and MME benchmarks.
@@ 6.2 Experimental Results
MSCOCO Random Regular $83.29_{(\pm 0.35)}$ $92.13_{(\pm 0.54)}$ $72.80_{(\pm 0.57)}$ $81.33_{(\pm 0.41)}$
VCD $87.73_{(\pm 0.40)}$ $91.42_{(\pm 0.55)}$ $72.80_{(\pm 0.42)}$ $87.16_{(\pm 0.41)}$
VDD $90.0_{(\pm 0.35)}$ $97.36_{(\pm 0.54)}$ $79.13_{(\pm 0.57)}$ $88.79_{(\pm 0.41)}$
Ours $\textbf{90.21}_{(\pm 0.50)}$ $\textbf{93.23}_{(\pm 0.55)}$ $\textbf{86.85}_{(\pm 0.42)}$ $\textbf{89.94}_{(\pm 0.41)}$
Popular Regular $81.88_{(\pm 0.48)}$ $88.93_{(\pm 0.60)}$ $72.80_{(\pm 0.57)}$ $80.06_{(\pm 0.05)}$
VCD $85.38_{(\pm 0.35)}$ $86.92_{(\pm 0.54)}$ $83.28(_{(\pm 0.57)}$ $85.06_{(\pm 0.41)}$
VDD $85.9_{(\pm 0.35)}$ $94.3_{(\pm 0.54)}$ $76.33_{(\pm 0.57)}$ $84.4_{(\pm 0.41)}$
Ours $\textbf{88.1}_{(\pm 0.40)}$ $88.96_{(\pm 0.55)}$ $\textbf{86.83}_{(\pm 0.42)}$ $\textbf{87.84}_{(\pm 0.41)}$
Adversarial Regular $78.96_{(\pm 0.52)}$ $83.06_{(\pm 0.58)}$ $72.75_{(\pm 0.59)}$ $77.57_{(\pm 0.57)}$
VCD $80.88_{(\pm 0.35)}$ $79.45_{(\pm 0.54)}$ $83.29_{(\pm 0.57)}$ $81.33_{(\pm 0.41)}$
VDD $83.5_{(\pm 0.35)}$ $89.3_{(\pm 0.54)}$ $76.2_{(\pm 0.57)}$ $82.2_{(\pm 0.41)}$
Ours $\textbf{84.32}_{(\pm 0.40)}$ $84.28_{(\pm 0.55)}$ $\textbf{84.33}_{(\pm 0.42)}$ $\textbf{84.34}_{(\pm 0.41)}$
@@
5 Truncate (length) 79.9 41.8 13.2 72.5
6 Truncate (proportion) 79.8 43.4 14.0 74.3
7 Truncate (sentence) 80.0 42.8 14.1 74.2
@@
LLaVA1.5 Regular $175.67_{(\pm 7.51)}$ $124.67_{(\pm 19.59)}$ $114.00_{(\pm 9.32)}$ $151.00_{(\pm 10.45)}$ $565.33_{(\pm 32.92)}$
VCD ${184.66}_{(\pm 6.81)}$ ${138.33}_{(\pm 15.68)}$ ${128.67}_{(\pm 7.21)}$ ${153.00}_{(\pm 7.58)}$ ${604.66}_{(\pm 18.76)}$
VDD ${190.0}_{(\pm 6.81)}$ ${143.3}_{(\pm 15.68)}$ ${145.0}_{(\pm 7.21)}$ ${165.0}_{(\pm 7.58)}$ ${643.3}_{(\pm 18.76)}$
Ours $\textbf{190.0}_{(\pm 6.81)}$ $\textbf{160.0}_{(\pm 15.68)}$ ${133.3}_{(\pm 7.21)}$ $\textbf{170.0}_{(\pm 7.58)}$ $\textbf{653.3}_{(\pm 18.76)}$
@@ 6.3 Ablation Study
Table 4: Ablation studies on different components of our method on MSCOCO Lin et al. (2014) dataset of LVLMs with different decoding baselines and SOTAs designed for mitigating object hallucination. Lower $\text{CHAIR}_{S}$ and $\text{CHAIR}_{I}$ indicate less object hallucination. Higher Recall generally represent higher recognition quality. Bold indicates the best results of all methods.
Appendix A Mathematical Derivations
$\sum_{i=k_{1}}^{k_{m}}({\hat{l_{i}}^{\{v,x,y_{<t}\}}}-{\hat{l_{j}}^{\{v,x,y_{<t}\}}})>J$
Appendix B Algorithm
This algorithm meticulously outlines the process enabling the model to autonomously generate a spectrum of potential hallucinations. Utilizing the pairs of samples thus produced, we employ our proposed Hallucination-Induced Optimization (HIO) to enhance the disparity between hallucination tokens and target tokens. Ultimately, the hallucinations are mitigated through the application of contrast decoding.
Algorithm 1 Training LVLM to Amplify Hallucination
@@ Appendix B Algorithm
Appendix C Additional experiments
[…]
Method $\text{CHAIR}_{S}\downarrow$ $\text{CHAIR}_{I}\downarrow$
Greedy $\text{20.80}_{\pm\text{0.08}}$ $\text{6.77}_{\pm\text{0.07}}$
Beam Search $\text{18.67}_{\pm\text{0.38}}$ $\text{6.30}_{\pm\text{0.05}}$
DoLA $\text{21.00}_{\pm\text{0.67}}$ $\text{6.70}_{\pm\text{0.38}}$
OPERA $\text{21.13}_{\pm\text{0.12}}$ $\text{6.73}_{\pm\text{0.18}}$
VCD $\text{23.33}_{\pm\text{5.66}}$ $\text{7.90}_{\pm\text{0.53}}$
Woodpecker $\text{23.85}_{\pm\text{4.62}}$ $\text{7.50}_{\pm\text{0.01}}$
LURE $\text{19.48}_{\pm\text{2.35}}$ $\text{6.5}_{\pm\text{0.38}}$
HALC ${13.80}_{\pm\text{0.08}}$ ${5.50}_{\pm\text{0.14}}$
Ours $\textbf{11.6}_{\pm\text{0.18}}$ $\textbf{2.24}_{\pm\text{0.26}}$
Table 5: CHAIR evaluation results on MSCOCO dataset of LVLMs with different decoding baselines and SOTAs designed for mitigating OH. Lower $\text{CHAIR}_{S}$ and $\text{CHAIR}_{I}$ indicate less object hallucination. Bold indicates the best results of all methods.
MME. To evaluate HIO’s capability on object-level and attribute-level hallucination, we compare it with several state-of-the-art Decoding methods on MME. The results are shown in Tab. 3. Consistent with the performance on POPE and CHAIR, HIO also achieves competitive results on MME compared to other decoding methods. Concretely, HIO outperforms the VCD at Existence, Count, Position, Color, Posters, on MME, respectively. The results demonstrate the effectiveness of our method.
@@ Appendix C Additional experiments
Total
LLaVA1.5 Regular $175.67_{\pm 7.51}$ $124.67_{\pm 19.59}$ $114.00_{\pm 9.32}$ $151.00_{\pm 10.45}$ $127.82_{\pm 7.13}$ $113.59_{\pm 3.43}$ $148.30_{\pm 3.49}$ $129.95_{\pm 5.33}$ $102.20_{\pm 4.70}$ $92.00_{\pm 31.29}$ $1279.19_{\pm 37.09}$
VCD ${184.66}_{\pm 6.81}$ ${138.33}_{\pm 15.68}$ ${128.67}_{\pm 7.21}$ ${153.00}_{\pm 7.58}$ ${132.11}_{\pm 6.53}$ ${120.94}_{\pm 7.57}$ ${152.20}_{\pm 0.21}$ ${140.45}_{\pm 6.73}$ ${109.60}_{\pm 2.66}$ ${104.00}_{\pm 30.96}$ ${1363.96}_{\pm 40.58}$
Ours $\textbf{190.0}_{\pm 6.81}$ $\textbf{160.0}_{\pm 15.68}$ $\textbf{133.3}_{\pm 7.21}$ $\textbf{170.0}_{\pm 7.58}$ $\textbf{145.5}_{\pm 6.53}$ $\textbf{138.5}_{\pm 7.57}$ $\textbf{158.7}_{\pm 0.21}$ $\textbf{165}_{\pm 6.73}$ $\textbf{121}_{\pm 2.66}$ $\textbf{142.5}_{\pm 30.96}$ $\textbf{1524.7}_{\pm 40.58}$

### v2

@@ 6.1 Experimental Settings
Benchmarks. We evaluate HIO on three benchmarks including: (1) Quantitative metrics POPE Li et al. (2023d) on MSCOCO Lin et al. (2014) dataset. The Polling-based Object Probing Evaluation Li et al. (2023d) offers a streamlined approach to assessing object hallucination. In this benchmark, LVLMs are queried about the existence of specific objects in a given image. (2) CHAIR Rohrbach et al. (2018), Caption Hallucination Assessment with Image Relevance, is a specialized tool designed to evaluate the occurrence of object hallucination in image captioning tasks. (3) General-purposed Multimodal Large Language Model Evaluation (MME) Fu et al. (2023) benchmark, which provides an extensive benchmark designed to evaluate LVLMs across multiple dimensions, including ten perception-related subtasks and four cognition-focused ones.
Implementation Description We evaluate our model across three Large Vision-Language Models (LVLMs): LLaVA 1.5, InstructBLIP, and MiniGPT-4. For decoding, we use Llama-7B and Vicuna-7B as the linguistic decoder for LLaVA and InstructBLIP/MiniGPT-4, respectively. Our model’s performance is compared against three leading models in the field: OPERA Huang et al. (2023), VCD Leng et al. (2023), and VDD Zhang et al. (2024). To ensure a fair and rigorous comparison, we adhere to the configurations and guidelines from the original works and codebases of the compared models. The training is conducted on a robust computational setup: 4x RTX 3090 GPUs for LLaVA 1.5, 8x V100 GPUs for MiniGPT-4, and 4x A6000 GPUs for InstructBLIP. Each training session lasts approximately 2-4 hours. Hyperparameters including alpha and beta are set to 1.0 and 0.1, respectively, in accordance with the VCD model’s specifications.
@@ 6.2 Experimental Results
MSCOCO Random Regular 83.29 92.13 72.80 81.33
VCD 87.73 91.42 72.80 87.16
ICD 89.56 88.71 90.66 89.68
VDD 90.00 97.36 79.13 88.79
Ours 90.21 93.23 86.85 89.94
Popular Regular 81.88 88.93 72.80 80.06
VCD 85.38 86.92 83.28 85.06
ICD 86.16 83.18 90.66 86.76
VDD 85.91 94.33 76.33 84.40
Ours 88.12 88.96 86.83 87.84
Adversarial Regular 78.96 83.06 72.75 77.57
VCD 80.88 79.45 83.29 81.33
ICD 79.71 74.35 90.66 81.70
VDD 83.52 89.34 76.20 82.20
Ours 84.32 84.28 84.33 84.34
@@
5 ICD 106.3 50.8 15.0 78.5
@@
LLaVA1.5 Regular $175.67$ $124.67$ $114.00$ $151.00$ $565.33$
VCD 184.66184.66{184.66}184.66 138.33138.33{138.33}138.33 128.67128.67{128.67}128.67 153.00153.00{153.00}153.00 604.66604.66{604.66}604.66
VDD 190.00190.00{190.00}190.00 143.33143.33{143.33}143.33 145.00145.00{145.00}145.00 165.00165.00{165.00}165.00 643.33643.33{643.33}643.33
Ours 190.00 160.00 133.33133.33{133.33}133.33 170.00 653.33
@@ 6.3 Ablation Study
Table 4: Ablation study with different components of our model on CHAIR-COCO.
Moreover, we have enriched the ablation study to analyze the generalization capability of our proposed components to unseen categories, as detailed in Table 4. For the Unseen-P dataset, we collected data from MSCOCO, A-OKVQA, and GQA, ensuring no overlap with the training set, resulting in 495 samples across 10 distinct classes. These experiments show that our components generalize effectively to unseen data. Finally, we have integrated the ablation study into the experimental results section, rather than presenting it separately.
[…]
- - - 88.88 84.88 95.63 83.93
- - 89.79 86.22 95.63 90.68
unseen-N - 91.83 95.30 88.64 91.85
92.97 91.94 94.75 93.33
- - - 81.15 64.86 100.00 78.68
- - 82.61 66.66 100.00 80.02
unseen-P - 84.05 72.41 87.51 79.24
85.51 75.01 87.51 80.76
Table 5: Ablation study on the generalization of each component on unseen datasets.
7 Discussion
Limitations &\&& Future Work.
8 Acknowledgments and Disclosure of Funding
This study is supported by grants from the National Natural Science Foundation of China (Grant No. 62122018, No. 62020106008, No. U22A2097, No. U23A20315), and Kuaishou.
Appendix
Appendix A Algorithm
The algorithm outlines the process by which the model generates its own series of potential hallucinations. Using the sample pairs produced by the model, we apply our proposed Hallucination-Induced Optimization (HIO) to enhance the distinction between hallucinated and target labels. Ultimately, hallucinations are mitigated through contrastive decoding.
Using paired hallucination and non-hallucination annotations from the RLHF-V dataset, we apply beam search to generate multiple outputs where hallucination token annotations occur. These outputs include both correct and hallucinated results, which we use as hallucination samples to reinforce the model’s confidence in its outputs. The correct annotations from RLHF-V serve as ground truth, helping the model avoid hallucinations by differentiating between hallucinated and target tokens. This approach expands the contrast between hallucinated and target tokens, effectively reducing hallucinations.
Algorithm 1 Training LVLM to Amplify Multiple Targeted Hallucination
@@ Appendix A Algorithm
Appendix B Mathematical Derivations
$\sum_{i=k_{1}}^{k_{m}}({\hat{l_{i}}^{\{v,x,y_{<t}\}}}-{\hat{l_{j}}^{\{v,x,y_{<t}\}}})>J$
Appendix C Visualization
This figure demonstrates the effectiveness of our ACI method (described in Section 4.3). The y-axis shows the difference between the hallucination token and the target token. The blue curve represents this difference without ACI, while the orange curve represents it with our proposed ACI. Clearly, our method accurately induces hallucinations, further amplifies the difference between the hallucination token and the target token, and thus effectively reduces hallucinations.
Appendix D Additional experiments
MME. To evaluate HIO’s capability on object-level and attribute-level hallucination, we compare it with several state-of-the-art Decoding methods on MME. The results are shown in Tab. 3. Consistent with the performance on POPE and CHAIR, HIO also achieves competitive results on MME compared to other decoding methods. Concretely, HIO outperforms the VCD at Existence, Count, Position, Color, Posters, on MME, respectively. The complete POPE evaluation is shown in the Tab 7.
@@ Appendix D Additional experiments
LLaVA1.5 Regular 175.67 $124.67$ $114.00$ $151.00$ $127.82$ $113.59$ $148.30$ $129.95$ $102.20$ $92.00$ $1279.19$
VCD 184.66184.66{184.66}184.66 138.33138.33{138.33}138.33 128.67128.67{128.67}128.67 153.00153.00{153.00}153.00 132.11132.11{132.11}132.11 120.94120.94{120.94}120.94 152.20152.20{152.20}152.20 140.45140.45{140.45}140.45 109.60109.60{109.60}109.60 104.00104.00{104.00}104.00 1363.961363.96{1363.96}1363.96
Ours 190.00 160.00 133.33 170.00 145.50 138.50 158.70 165.00 121.00 142.50 1524.70